Consider a perpetuity whose payments at the end of each year are R, R+p, R+2p,..., R+(n-1)p, R+np, R+np,... . The payments increase by a constant amount p until they reach R+np, after which they continue without change. Show that the discounted value A of such a perpetuity at rate i per annum is given by:


A = (R+pa(n)i) / i


1
Expert's answer
2021-11-04T08:42:46-0400

The formula for determining A is given by:


A="\frac{R}{(1+i)^1}" +"\frac{R}{(1+i)^2}" +"\frac{R}{(1+i)^3}" ......+"\frac{R}{(1+i)^n}" ="\sum^\infty_{n=1}" "\frac{R}{(1+i)^n}" ="\frac{R}{i}" ................equation 1


Where R = the payment or receipt each period

i = the interest rate per payment or receipt period 


The discount at a perpetuity rate i per annum is calculated by;


"pa(n)^{i}" ......................................................equation 2


Where pa= the amount

i = the interest rate per payment or receipt period


n=number of years


Inserting equation 2 into value of R in equation 1 to get the new value of R at a discounted rate, we get the discounted value A as:


A = "\frac{(R+pa(n)^{i})}{i}"